calculate the pH: 35ml 0.2M formic acid 55mL of 0.1M sodium formate and 110mL of DI Ka of formic acid is 1.8 x 10 -4
Answer #1

Ka of formic acid, HCOOH = 1.8810-4 þka = -log ka = -log (1.8810-4) - 3.745 moles = volume (L x concentration (M) (1000ml = 1